OpenStudy (anonymous):

Solve for z. - 9 - 6z = - 99 plz help

OpenStudy (mikurout):

-9-6z +9 = -99 +9 -6z = -99 -(-6z) = -(-99) 6z = 99 (6z) /6 = 99/6 z = 99/6

OpenStudy (mikurout):

sorry a little mistake in 2nd line.

OpenStudy (luigi0210):

Please don't do all the work for them >.< Btw you messed up in step two :P

OpenStudy (mikurout):

it is -90

OpenStudy (anonymous):

can you explain it for me

OpenStudy (anonymous):

so i know how to do it

OpenStudy (mikurout):

Where is your doubt?

OpenStudy (anonymous):

what does the z do

OpenStudy (mikurout):

It is just a variable.

OpenStudy (anonymous):

well my question is how did you get -90

OpenStudy (anonymous):

im kinda new to this

OpenStudy (mikurout):

-9 - 6z = -99 for getting the value of z 1st remove that (-9) from left hand side. to do this add +9 to both side.

OpenStudy (anonymous):

ok

OpenStudy (mikurout):

In a equation you have to do the same operation in both side . otherwise it will not remain as a equation.

OpenStudy (mikurout):

So in LHS after adding +9 , you will get -6z because that +9 and -9 will give you 0 and 0 -6z = -6z

OpenStudy (anonymous):

oh

OpenStudy (mikurout):

In RHS you have -99 . you are adding +9 to it. so -99 +9 will give you -90 To understand this, just do the substraction from greater number to smaller one. this will give you 99 - 9 = 90 then put the sign in your answer. sign of answer will be the sign of the greatest number. so here your answer will be - 90

OpenStudy (anonymous):

oh ok thank you so much

OpenStudy (mikurout):

@Luigi0210 Yes , it should be -90.